# Modeling information spread across networks with communities using a multitype branching process framework

### v1

@@ Abstract
The dynamics of information diffusion in complex networks is widely studied in an attempt to understand how individuals communicate and how information travels and reaches individuals through interactions. However, complex networks often present community structure, and tools to analyse information diffusion on networks with communities are needed. In this paper, we develop theoretical tools using multi-type branching processes to model and analyse simple contagion information spread on a broad class of networks with community structure. We show how, by using limited information about the network — the degree distribution within and between communities — we can calculate standard statistical characteristics of the dynamics of information diffusion, such as the extinction probability, hazard function, and cascade size distribution. These properties can be estimated not only for the entire network but also for each community separately. Furthermore, we estimate the probability of information spreading from one community to another where it is not currently spreading. We demonstrate the accuracy of our framework by applying it to two specific examples: the Stochastic Block Model and a log-normal network with community structure. We show how the initial seeding location affects the observed cascade size distribution on a heavy-tailed network and that our framework accurately captures this effect.
@@ I Introduction and Background
Understanding the dynamics of information spread across complex networks is crucial in today’s interconnected world amedie_impact_2015 . Platforms such as Facebook and X (formerly Twitter) which connect millions of users and facilitate information spreading among them, have a profound impact on human societies. Social media shapes public opinion on personal as well as societal issues, including public health clark_role_2022 ; johnson_online_2020 ; noauthor_infodemics_nodate ; beguerisse-diaz_who_2017 , politics amedie_impact_2015 ; persily_social_2020 , and business aral_distinguishing_2009 . It is, therefore, crucial to understand how information propagates among agents connected via complex networks and what factors accelerate or slow down the spread. Consequently, many efforts have been devoted to developing mathematical tools to analyse propagation on networks boguna_absence_2003 ; miller_percolation_2009 ; porter_dynamical_2016 . Information diffusion shares many characteristics with epidemic processes. However, in the study of information diffusion, we almost always deal with the process unfolding on a complex network of connections, where the structure of the network becomes crucial in determining the dynamics. These networks can be complex, but there are ways to measure and summarise them, with many characteristics being well-known.
@@
Community structure is one of the most discussed structural properties of social networks. Previous studies have highlighted the importance of community structure roles in explaining the spread of social contagions via reinforcement and homophily using both empirical and statistical analyses weng_virality_2013 ; zhou_survey_2022 . Some have used communities as a feature to predict cascade size using machine learning and statistical methods hoang_gpop_2017 ; weng_predicting_2014 . Community structure in real-world networks has been shown to have a substantial impact on percolation and epidemic spreading processes, where the community structure can both benefit and slow down the diffusion processes stegehuis_epidemic_2016 . Additionally, mathematical models for spreading processes have also been developed, where probability-generating functions (PGFs) are used to capture the network and community structure, and then standard ordinary differential equations compartmental models are applied to model the diffusion processes li_sir_2018 . However, the literature is scarce when it comes to theoretical studies of information spread on a network with community structure using branching processes li_influence_2018 ; gomez-rodriguez_inferring_2012 ; peixoto_network_2019 . One notable exception is Brummitt et al.’s brummitt_suppressing_2012 work on the modelling of interconnected electricity grids, which showed the impact of sparsely connected electricity networks on the occurrence of blackouts of various sizes.
@@
The branching processes framework allows us to estimate the distributional properties of the process. Instead of focusing on the average behaviour of cascades, which has been the main focus of many previous works, in this paper, we are interested in estimating distributions of the typical statistical characteristics of the cascade dynamics, e.g., cascade size distribution and extinction time distribution. Branching process approximation has previously proven effective in studying information diffusion across local tree-like networks gleeson_branching_2021 , as well as on both homogeneous clique type networks and networks with heterogeneous degree and clustering keating_multitype_2022 ; Keating_Gleeson_OSullivan_2023 . These works laid down a theoretical formalism that we can extend to networks with community structure. The introduction of community structure brings extra complexity to the analysis of information cascades. Information diffusion may halt within a community but be reintroduced later from another community, which, for example, will affect the calculation of the probability that information stops spreading (extinction probability) in a community. This complexity requires the development of new mathematical methods to describe information diffusion across communities.
@@
In this paper, we develop theoretical tools to model and analyse information spread on a broad class of networks with community structure for a simple contagion mechanism that follows the Independent Cascade Model (ICM) kempe_maximizing_2003 ; lerman_information_2016 . In the ICM model, each exposure of a susceptible individual to a disease (or piece of information) by an infected friend (or followed person) results in an independent chance of disease transmission (information adoption). We use a multi-type branching process approximation caswell_matrix_2002 and probability-generating functions to describe the spread under the ICM. We derive analytical expressions for fundamental characteristics of information spread, such as the probability of extinction, hazard function and cascade sizes distribution. Furthermore, we derive specific features associated with cross-community diffusion, such as the introduction probability — the probability of information (re)introduction to a community where it is not currently spreading. We demonstrate the effectiveness of our analysis initially on the Stochastic Block Model and then extend it to a general locally tree-like network with communities with known degree distributions.
@@
II Conceptual model description for information spread on networks with communities
In this paper, we study information diffusion following the Independent Cascade Model on community-based networks. The schematic of the process along with the type of network we study is illustrated in Fig. 1a). The Independent Cascade Model (ICM) defines nodes of a network as being in one of three states: inactive, active, and removed. Active nodes independently attempt to activate their inactive neighbors with probability $\rho$, transitioning to the removed state afterward. Once the node is in the removed state, it cannot be reactivated. The process begins with a randomly selected active node from community 1 and terminates when no new active nodes are generated. We consider information diffusion that can be modelled as a subcritical branching process, which is certain to terminate and only spread to a fraction of the possible nodes in a network. Such processes are interesting as many of the social processes that we wish to capture are subcritical, where, even the largest diffusions only reach a small fraction of the potential audience, e.g., even the largest Twitter diffusion only reaches a small fraction of users gleeson_branching_2021 .
Refer to captiona)
Refer to captionb)
Figure 1: a) Schematic illustration of the Independent Cascade Model (ICM) on a community-based network. The network consists of two communities each described by its internal degree distributions and distribution of edges between communities. Here $D_{1}^{(1)}$ and $D_{2}^{(2)}$ are random variables for the nodes degree inside community 1 and 2 respectively; $D_{2}^{(1)}$ and $D_{1}^{(2)}$ are random variables for a nodes degree between-community (they are the same in the case of undirected networks considered here). At each time step, active nodes attempt to activate neighbours with probability $\rho$, then become inactive (or “removed”). The information spreads in the first community initially and later propagates to the second community. b) Schematic of the multi-type branching process approximating ICM on a community-based network. We track two types of offspring: “type 1” for active nodes in community 1, and “type 2” for active nodes in community 2. Here $\boldsymbol{N}(t)=(N_{1}(t),N_{2}(t))$ tracks the number of active nodes of type 1 ($N_{1}(t)$) and type 2 ($N_{2}(t)$) at each time step.
@@ II Conceptual model description for information spread on networks with communities
Our goal is to calculate typical statistical characteristics of information diffusion, such as total cascade size distribution (also referred to as the total outbreak size or total progeny distribution) and extinction probability. Additionally, we aim to extract quantities that are particularly interesting for information diffusion on community-based networks, such as the probability of a contagion “escaping” its seed community or the probability of reintroduction of a contagion into a community where it has previously stopped spreading.
To calculate these quantities, we will describe the information diffusion process through a multi-type branching process. A multi-type branching process (MTBP) is a discrete Markov process that tracks the active nodes from each community at each time step as random variables caswell_matrix_2002 ; Keating_Gleeson_OSullivan_2023 ; kimmel_branching_2002 . We will refer to the active nodes in community 1 as “type 1” active nodes, and the active nodes in community 2 will be called “type 2” active nodes. At each time step, also referred to as a generation of the process, each type of node independently produces a random number of active children node. These active children are refereed to as the active parents offspring in the Branching Processes literature caswell_matrix_2002 . The offspring follow a specific probability distribution, known as the offspring distribution. These offspring then continue the process, generating new descendants in subsequent generations, as illustrated in Fig. 1b). The primary components in this formulation are the offspring distributions, one specified for each type of node that we track. These distributions are then used to determine the number of active nodes of each type in a given generation (time step). The goal, generally in a Branching Processes, is to calculate the probability distribution of the total number of offspring of each type produced in each generation, as this fully determines the dynamics of the branching process caswell_matrix_2002 . Once this distribution is obtained, we can proceed with calculating key characteristics, such as cascade size and extinction probabilities, etc.
@@ III Probability-generating function framework in a Poisson-distributed network
Now let us calculate the offspring distribution produced by a node of community 1, but note the same logic will apply to the nodes of the second community. First, we calculate the number of activation a node of community 1 produces inside its community. The probability of a node having $k$ neighbours inside community 1 is given by Eqn. (5). Each of these neighbour nodes is activated independently with probability $\rho$, meaning that the probability of having $x$ successfully activated nodes in community 1 from an active node with in-degree $k$ is $\dfrac{\lambda_{in}^{k}e^{\lambda_{in}}}{k!}{k\choose x}(1-\rho)^{k-x}(\rho)^{x}$. Summing over all possible values of $k$ yields the probability of having $x$ activated offspring nodes inside community 1 as follows
@@
We have the probability generating function for the number of active nodes of type 1 and type 2 at time $t-1$. Using this, we wish to see how many offspring we will have in the following generation using $\boldsymbol{X^{(1)}}$ and $\boldsymbol{X^{(2)}}$. This characterisation of the branching process is called the forward approach in Ref. kimmel_branching_2002 , in analogy with the forward Chapman-Kolmogorov equation of Markov processes. It is the common way in which one can derive the number of active nodes in a generation once we have obtained the offspring distributions for each type in MTBP. Refer to Caswell caswell_matrix_2002 for a derivation of this relationship. We set the initial condition $G_{\boldsymbol{N}(0)}(s_{1},s_{2})=(s_{1})^{1}(s_{2})^{0}$ corresponding to a single active individual in community 1 and none in community 2.
@@
To obtain the probability distribution of number of active nodes in each of the communities separately, $P[N_{1}(t)=n]$ and $P[N_{2}(t)=m]$, we simply marginalise the probabilities in the community that is not of interest. Thus, the probability $P[N_{1}(t)=n]$ will be recovered from $G_{\boldsymbol{N}(t)}(s_{1},1)$, while the probability $P[N_{2}(t)=m]$ will be recovered from $G_{\boldsymbol{N}(t)}(1,s_{2})$, e.g.,
@@ III.1 Extinction probability and hazard function
With a bit more effort, we can compute the hazard function, which is the probability that the process is extinct in a community at time $t$ given that it is not extinct at $t-1$, while possibly still spreading in the other community. We start with the hazard function for community 1, $h_{1}(t)=P[N_{1}(t)=0|N_{1}(t-1)+N_{2}(t-1)>0]$. This can be expressed as
@@ III.1.1 Community specific hazard function
We can now calculate what we call the community specific hazard function, $\widetilde{h}_{1}(t)$ and $\widetilde{h}_{2}(t)$. They are hazard functions specified for each community on non-extinction in community 1 or community 2 in the previous time step, i.e., $\widetilde{h}_{1}(t)=P[N_{1}(t)=0|N_{1}(t-1)>0]$ and $\widetilde{h}_{2}(t)=P[N_{2}(t)=0|N_{2}(t-1)>0]$. In contrast, $h_{1}(t)$ in the last section was the hazard function for community 1 conditioned on the joint non-extinction of the process in the previous time step (i.e., $N_{1}(t-1)=0$ and $N_{2}(t-1)$). This can be useful if we care more about when the process stops in a specific community disregarding whether it is still going on in the other one. To derive the expression for $\widetilde{h}_{2}(t)$ we introduce the random variables $Y_{1}(t)$ and $Y_{2}(t))$ such that $P[Y_{1}(t)=k]=P[N_{1}(t)=k|N_{1}(t-1)>0]$. We note that $Y_{1}(s)$ is the PGF for the number of active nodes at time $t$, conditional on the process being non-zero in community 1. Then $G_{Y_{1}(t)}(s_{1},s_{2})$ is easily calculated from $G_{\boldsymbol{N}(t)}(s_{1},s_{2})$, as follows
@@
To see this, write out the power series of $G_{\boldsymbol{N}(t)}(s_{1},s_{2})$ and subtract the terms corresponding to $G_{\boldsymbol{N}(t)}(0,s_{2})$, which then is normalised. We can now note a connection between our desired probability $\widetilde{h}_{1}(t)$ and $G_{Y_{1}(t-1)}$. The function $G_{Y_{1}(t-1)}$ tracks the probabilities for the number of active nodes at time $t-1$ given the process is not extinct in community 1. If we find the number of offspring of the active nodes from the $t-1$ generation, we can arrive at the PGF for number of active nodes in generation $t$ conditioned on the process not being extinct in generation $t-1$. This is given by:
@@
This gives us a way to express the community-specific hazard probability for community 1. Then by calculating the probability of extinction in community 1 (setting $s_{1}=0$) and marginalising over community 2 (by setting $s_{2}=1$) in generation $t$, we obtain $\widetilde{h}_{1}(t)$, as required:
@@
Similarly, the community-specific hazard probabilities for community 2, $\widetilde{h}_{2}(t)$, can be derived through an analogous process. It is important to note that we can arrive at all these expressions via simple function iteration followed by the appropriate substitutions of zeros and ones into the dummy variables $s_{1}$ and $s_{2}$. Equation (21) allows us to finally arrive at an important quantity of interest, the probability of information spreading between communities, which we derive in the next section.
III.2 Probability of information travel between communities
When studying the spread of information or disease on a network with communities, a crucial question arises: What is the probability of infection spreading to another community? Here we show how to calculate this probability with our framework. Let $r_{1}(t)=P[N_{1}(t)>0|N_{1}(t-1)=0]$ be the probability of infection in the first community at time $t$, given no infected nodes in that community at time $t-1$, which we will call the introduction probability for community 1. Care must be taken in noting what $r_{1}(t)$ is; it measures the probability of seeing an activated node in community 1, maybe for the first time, but also for the second time, third time, etc. All we know is that there was no infected (active) node in the community in the previous time step. It can be expressed as

### v2

# Modeling diffusion in networks with communities: a multitype branching process approach

@@ Abstract
The dynamics of diffusion in complex networks are widely studied to understand how entities, such as information, diseases, or behaviors, spread in an interconnected environment. Complex networks often present community structure, and tools to analyze diffusion processes on networks with communities are needed. In this paper, we develop theoretical tools using multi-type branching processes to model and analyze diffusion processes, following a simple contagion mechanism, across a broad class of networks with community structure. We show how, by using limited information about the network — the degree distribution within and between communities — we can calculate standard statistical characteristics of propagation dynamics, such as the extinction probability, hazard function, and cascade size distribution. These properties can be estimated not only for the entire network but also for each community separately. Furthermore, we estimate the probability of spread crossing from one community to another where it is not currently spreading. We demonstrate the accuracy of our framework by applying it to two specific examples: the Stochastic Block Model and a log-normal network with community structure. We show how the initial seeding location affects the observed cascade size distribution on a heavy-tailed network and that our framework accurately captures this effect.
@@ I Introduction and Background
Understanding the dynamics of spreading phenomena across complex networks is crucial in today’s interconnected world siddiqui2016social. Various forms of spread, whether they involve diseases, innovations, behaviors, or information, have a profound impact on human societies. The way diseases spread can determine public health outcomes meagher_assessing_2022; fyles_using_2021, the way innovations diffuse can shape economic development mccullen2013multiparameter; peres2014impact, and the way behaviors and ideas propagate can influence social and political landscapes siddiqui2016social; persily_social_2020; juul2019hipsters; aral_distinguishing_2009; clark_role_2022; johnson_online_2020; noauthor_infodemics_nodate; beguerisse-diaz_who_2017. It is, therefore, crucial to understand propagation dynamics among agents connected via complex networks and what factors accelerate or slow down the spread. Consequently, many efforts have been devoted to developing mathematical tools to analyze propagation on networks boguna_absence_2003; miller_percolation_2009; porter_dynamical_2016. In the study of diffusion processes, we often deal with the process unfolding on a complex network of connections, where the structure of the network becomes crucial in determining the dynamics. These networks can be complex, but there are ways to measure and summarize them, with many characteristics being well-known.
@@
Community structure is one of the most discussed structural properties of social and physical contact networks. Previous studies have highlighted the importance of community structure roles in explaining the spread of social contagions via reinforcement and homophily using both empirical and statistical analyses weng_virality_2013; zhou_survey_2022. Some have used communities as a feature to predict cascade size using machine learning and statistical methods hoang_gpop_2017; weng_predicting_2014. Community structure in real-world networks has been shown to have a substantial impact on percolation and epidemic spreading processes, where the community structure can both benefit and slow down the diffusion processes stegehuis_epidemic_2016. Additionally, mathematical models for spreading processes have also been developed, where probability-generating functions (PGFs) are used to capture the network and community structure, and then standard ordinary differential equations compartmental models are applied to model the diffusion processes li_sir_2018. However, the literature is scarce when it comes to theoretical studies of propagation dynamics on a network with community structure using branching processes li_influence_2018; gomez-rodriguez_inferring_2012; peixoto_network_2019. One notable exception is Brummitt et al.’s brummitt_suppressing_2012 work on modeling interconnected electricity grids, which showed the impact of sparsely connected electricity networks on the occurrence of blackouts of various sizes.
@@
The branching processes framework allows us to estimate the distributional properties of the process. Instead of focusing on the average behavior of cascades, which has been the main focus of many previous works, in this paper, we are interested in estimating distributions of the typical statistical characteristics of the cascade dynamics, e.g., cascade size distribution and extinction time distribution. Branching process approximation has previously proven effective in studying diffusion across local tree-like networks gleeson_branching_2021, as well as on both homogeneous clique type networks and networks with heterogeneous degree and clustering keating_multitype_2022; Keating_Gleeson_OSullivan_2023. These works laid down a theoretical formalism that we can extend to networks with community structure. The introduction of a community structure adds additional complexity to the analysis of cascades. Spreading may halt within a community but can later be reintroduced from another community, which, for example, affects the calculation of the probability that diffusion stops (extinction probability) within a community. This complexity requires the development of new mathematical methods to describe diffusion processes across communities.
@@
In this paper, we develop theoretical tools to model and analyze propagation dynamics on a broad class of networks with community structure for a simple contagion mechanism that follows the Independent Cascade Model (ICM) kempe_maximizing_2003; lerman_information_2016. In the ICM model, each exposure of a susceptible individual to a disease (or piece of information) by an infected friend (or followed person) results in an independent chance of disease transmission (information adoption). We use a multi-type branching process approximation caswell_matrix_2002 and probability-generating functions to describe the spread under the ICM. We derive analytical expressions for fundamental characteristics of spreading phenomena, such as the probability of extinction, hazard function, and cascade sizes distribution. Furthermore, we derive specific features associated with cross-community diffusion, such as the introduction probability — the probability of information (or disease) (re)introduction to a community where it is not currently spreading. We demonstrate the effectiveness of our analysis initially on the Stochastic Block Model and then extend it to a general locally tree-like network with communities with known degree distributions.
@@
II Conceptual model description of the spreading process on networks with communities
In this paper, we study the diffusion process following the Independent Cascade Model on community-based networks. The schematic of the process along with the type of network we study is illustrated in Fig. 1a). The Independent Cascade Model (ICM) defines nodes of a network as being in one of three states: inactive, active, and removed. Active nodes independently attempt to activate their inactive neighbors with probability $\rho$, transitioning to the removed state afterward. Once the node is in the removed state, it cannot be reactivated. The process begins with a randomly selected active node from community 1 and terminates when no new active nodes are generated. We consider diffusion that can be modeled as a subcritical branching process, which is certain to terminate and only spread to a fraction of the possible nodes in a network. Such processes are interesting as many of the social or disease spreading processes we wish to capture are subcritical, where, even the largest diffusion only reach a small fraction of the potential audience, e.g., even the largest Twitter diffusion only reaches a small fraction of users gleeson_branching_2021.
Figure 1: a) Schematic illustration of the Independent Cascade Model (ICM) on a community-based network. The network consists of two communities each described by its internal degree distributions and distribution of edges between communities. Here $D_{1}^{(1)}$ and $D_{2}^{(2)}$ are random variables for the nodes degree inside community 1 and 2 respectively; $D_{2}^{(1)}$ and $D_{1}^{(2)}$ are random variables for a nodes’ degree between-community (they are the same in the case of undirected networks considered here). At each time step, active nodes attempt to activate neighbors with probability $\rho$, then become inactive (or “removed”). The spread starts in the first community initially and later propagates to the second community. b) Schematic of the multi-type branching process approximating ICM on a community-based network. We track two types of offspring: “type 1” for active nodes in community 1, and “type 2” for active nodes in community 2. Here $\boldsymbol{N}(t)=(N_{1}(t),N_{2}(t))$ tracks the number of active nodes of type 1 ($N_{1}(t)$) and type 2 ($N_{2}(t)$) at each time step.
@@ II Conceptual model description of the spreading process on networks with communities
Our goal is to calculate typical statistical characteristics of the diffusion process, such as total cascade size distribution (also referred to as the total outbreak size or total progeny distribution) and extinction probability. Additionally, we aim to extract quantities that are particularly interesting for diffusion on community-based networks, such as the probability of a contagion “escaping” its seed community or the probability of reintroduction of a contagion into a community where it has previously stopped spreading.
To calculate these quantities, we will describe the diffusion process through a multi-type branching process. A multi-type branching process (MTBP) is a discrete Markov process that tracks the active nodes from each community at each time step as random variables caswell_matrix_2002; Keating_Gleeson_OSullivan_2023; kimmel_branching_2002. We will refer to the active nodes in community 1 as “type 1” active nodes, and the active nodes in community 2 will be called “type 2” active nodes. At each time step, also referred to as a generation of the process, each type of node independently produces a random number of active children node. These active children are refereed to as the active parent’s offspring in the Branching Processes literature caswell_matrix_2002. The offspring follow a specific probability distribution, known as the offspring distribution. These offspring then continue the process, generating new descendants in subsequent generations, as illustrated in Fig. 1b). The primary components in this formulation are the offspring distributions, one specified for each type of node that we track. These distributions are then used to determine the number of active nodes of each type in a given generation (time step). The goal, generally in a Branching Processes, is to calculate the probability distribution of the total number of offspring of each type produced in each generation, as this fully determines the dynamics of the branching process caswell_matrix_2002. Once this distribution is obtained, we can proceed with calculating key characteristics, such as cascade size, extinction probabilities, etc.
@@ III Probability-generating function framework in a Poisson-distributed network
Now let us calculate the offspring distribution produced by a node of community 1, but note the same logic will apply to the nodes of the second community. First, we calculate the number of activation a node of community 1 produces inside its community. The probability of a node having $k$ neighbors inside community 1 is given by Eqn. (5). Each of these neighbor nodes is activated independently with probability $\rho$, meaning that the probability of having $x$ successfully activated nodes in community 1 from an active node with in-degree $k$ is $\dfrac{\lambda_{in}^{k}e^{\lambda_{in}}}{k!}{k\choose x}(1-\rho)^{k-x}(\rho)^{x}$. Summing over all possible values of $k$ yields the probability of having $x$ activated offspring nodes inside community 1 as follows
@@
We have the probability generating function for the number of active nodes of type 1 and type 2 at time t1t-1. Using this, we wish to see how many offspring we will have in the following generation using $\boldsymbol{X^{(1)}}$ and $\boldsymbol{X^{(2)}}$. This characterization of the branching process is called the forward approach in Ref. kimmel_branching_2002, in analogy with the forward Chapman-Kolmogorov equation of Markov processes. It is the common way in which one can derive the number of active nodes in a generation once we have obtained the offspring distributions for each type in MTBP. Refer to Caswell caswell_matrix_2002 for a derivation of this relationship. We set the initial condition $G_{\boldsymbol{N}(0)}(s_{1},s_{2})=(s_{1})^{1}(s_{2})^{0}$ corresponding to a single active individual in community 1 and none in community 2.
@@
To obtain the probability distribution of the number of active nodes in each of the communities separately, $P[N_{1}(t)=n]$ and $P[N_{2}(t)=m]$, we simply marginalize the probabilities in the community that is not of interest. Thus, the probability $P[N_{1}(t)=n]$ will be recovered from $G_{\boldsymbol{N}(t)}(s_{1},1)$, while the probability $P[N_{2}(t)=m]$ will be recovered from $G_{\boldsymbol{N}(t)}(1,s_{2})$, e.g.,
@@ III.1 Extinction probability and hazard function
With a bit more effort, we can compute the hazard function, which is the probability that the process is extinct in a community at time $t$ given that it is not extinct at t1t-1, while possibly still spreading in the other community. We start with the hazard function for community 1, $h_{1}(t)=P[N_{1}(t)=0|N_{1}(t-1)+N_{2}(t-1)>0]$. This can be expressed as
@@ III.1.1 Community specific hazard function
We can now calculate what we call the community specific hazard function, $\widetilde{h}_{1}(t)$ and $\widetilde{h}_{2}(t)$. They are hazard functions specified for each community on non-extinction in community 1 or community 2 in the previous time step, i.e., $\widetilde{h}_{1}(t)=P[N_{1}(t)=0|N_{1}(t-1)>0]$ and $\widetilde{h}_{2}(t)=P[N_{2}(t)=0|N_{2}(t-1)>0]$. In contrast, $h_{1}(t)$ in the last section was the hazard function for community 1 conditioned on the joint non-extinction of the process in the previous time step (i.e., $N_{1}(t-1)=0$ and $N_{2}(t-1)$). This can be useful if we care more about when the process stops in a specific community disregarding whether it is still going on in the other one. To derive the expression for $\widetilde{h}_{1}(t)$ we introduce the random variable $Y_{1}(t)$ such that $P[Y_{1}(t)=k]=P[N_{1}(t)=k|N_{1}(t-1)>0]$. Then $G_{Y_{1}(t)}$ is the PGF for the number of active nodes at time $t$, conditional on the process being non-zero in community 1, and it is easily calculated from $G_{\boldsymbol{N}(t)}$, as follows
@@
To see this, write out the power series of $G_{\boldsymbol{N}(t)}(s_{1},s_{2})$ and subtract the terms corresponding to $G_{\boldsymbol{N}(t)}(0,s_{2})$, which then is normalized. We can now note a connection between our desired probability $\widetilde{h}_{1}(t)$ and $G_{Y_{1}(t-1)}$. The function $G_{Y_{1}(t-1)}$ tracks the probabilities for the number of active nodes at time t1t-1 given the process is not extinct in community 1. If we find the number of offspring of the active nodes from the t1t-1 generation, we can arrive at the PGF for number of active nodes in generation $t$ conditioned on the process not being extinct in generation t1t-1. This is given by:
@@
This gives us a way to express the community-specific hazard probability for community 1. Then by calculating the probability of extinction in community 1 (setting $s_{1}=0$) and marginalizing over community 2 (by setting $s_{2}=1$) in generation $t$, we obtain $\widetilde{h}_{1}(t)$, as required:
@@
Similarly, the community-specific hazard probabilities for community 2, $\widetilde{h}_{2}(t)$, can be derived through an analogous process. It is important to note that we can arrive at all these expressions via simple function iteration followed by the appropriate substitutions of zeros and ones into the dummy variables $s_{1}$ and $s_{2}$. Equation (21) allows us to finally arrive at an important quantity of interest, the probability of contagion spreading between communities, which we derive in the next section.
III.2 Probability of contagion travel between communities
When studying the spread of information or disease on a network with communities, a crucial question arises: What is the probability of infection spreading to another community? Here we show how to calculate this probability with our framework. Let $r_{1}(t)=P[N_{1}(t)>0|N_{1}(t-1)=0]$ be the probability of infection in the first community at time $t$, given no infected nodes in that community at time t1t-1, which we will call the introduction probability for community 1. Care must be taken in noting what $r_{1}(t)$ is; it measures the probability of seeing an activated node in community 1, maybe for the first time, but also for the second time, third time, etc. All we know is that there was no infected (active) node in the community in the previous time step. It can be expressed as